Pam runs a mail order business for gym equipment. Annual demand is 16,000. The cost of the item is $25. The annual holding cost per unit is 10% of the value of the item. The cost to place an order is $50. Find the optimal inventory level, EOQ. Then find the total cost (storage + order cost).
Answer #1

holding cost = 10% * 25 = 2.5

EOQ = [2 * Demand * Cost per order / holing cost] ^ (1/2)

= [2 * 16000 * 50 / 2.5] ^ (1/2)

= 640000 ^ 0.5

= 800

Total cost = (800 / 2) * 2.5 + (16000/800 )* 50 = 400 * 2.5 + 20 * 50 = 2000